which of the following points lie on the graph of the function X+3y=-7
Answer by arunpaul(104)

X+3y=-7
let x=1 then putting the value in the above equation we have
1+3y=-7
or 3y=-8
or y=-8/3
so the point (1,-8/3) lies in the line asked